OpenStudy (anonymous):

DIVISION QUESTION BELOW PLEASE HELP! (:

OpenStudy (anonymous):

Throw it at us

OpenStudy (anonymous):

heheh, lets play some catch with it

OpenStudy (anonymous):

\[26x \div -3\] this was part of a bigger problem but I solved it up until here although I'm not sure where to go from here.

OpenStudy (anonymous):

\[\frac{ 26x }{ -3 }\]that's as simply as the problem gets actually, if you came to the end with that, then your answer is as simple as i'ts going to get, unless your answer needs to fit some kind of form? Otherwise that's right, if you want to make sure, post up the full problem

OpenStudy (anonymous):

Was it an equation to find the value for x?

OpenStudy (anonymous):

No, the equation just said to divide, I probably did it all wrong though, math's not my strong subject, here's the full equation \[26xy - 39y \div 13y\]

OpenStudy (anonymous):

If your entire question looks like this\[\frac{ 26xy-39y }{ 13y }\]you have to remember that you're dividing both terms by the 13y, not just one of them, so, in other words your equation looks like this a well\[\frac{ 26xy }{ 13y }-\frac{ 39y }{ 13y }\]and this can be simplified very easily to be\[2x-3\]now when you write out the question using the division sign, you should add parenthesis like this\[(26xy-39y)\div13y\]otherwise you will always get the wrong answer if you follow the order of operations

OpenStudy (anonymous):

Oh okay that makes more sense, I think I figured out where I went wrong, thank you! (: